OpenStudy (anonymous):

Linear Transformation: Let T:R^2 ==> R^3. We know T(1,1)=(1,2,3) and T(1,2)=(3,4,5). Use linearity to compute. T(3,3)

OpenStudy (anonymous):

So, we know the functions are linear, thus: \[ z=a_1x+b_1y\\ w=a_2x+b_2y\\ v=a_3x+b_3y \]Given that: \[ z=1=a_1(1)+b_1(1)\\ z=3=a_1(1)+b_1(2)\implies\\ b_1=2\\ a_1=-2 \]Et al. Try that. Unless, by linearity of transforms you mean, where \(a, b\in\mathbb{R}^2\): \[ T(a)+T(b)=T(a+b) \]Then we may have multiple possible functions, I think.